Francois rolls a pair of dice.
A.) what is the probability that he will roll a total of 9 on one roll?

B.) what is the probability that he will roll a 5 followed by a 2 on two consecutive rolls?

Answer by rothauserc(4718)

A) using a pair of dice, the possible nines are
(3,6), (6,3) (4, 5), (5, 4)
There are 36 possible outcomes, therefore
************************************************
Probability of rolling a total of 9 = 4/36 = 1/9
************************************************
:
B) using a pair of dice,
(5,2), (5,2)
Probability of rolling (5,2) on two consecutive rolls is (1/36) * (1/36) = 1/1296
*********************************************************************************
: